OpenStudy (anonymous):

Your car averages 380 miles on a 14-gallon tank of gas. You run out of gas on a deserted highway, but have 4072-01-01-05-00_files/i0300000.jpg gallon of lawn mower gas with you. At the current rate of travel, how far will you be able to go and will you reach the nearest gas station 15 miles away? a. 25.33 miles; yes b. 7.0 miles; no c. 27.14 miles; yes d. 13.57 miles; no

OpenStudy (vianne):

can you resend photo? :P

OpenStudy (anonymous):

what pict

OpenStudy (anonymous):

@Vianne

OpenStudy (vianne):

"4072-01-01-05-00_files/i0300000.jpg" This isn't some HTML website where you can reply and attach an image using HTML Form :PP There is a button called "Attach File" (If you can see it) and can you just attach the file ? "4072-01-01-05-00_files/i0300000.jpg" is what it apparently is trying to be shown.

OpenStudy (anonymous):

Your car averages 380 miles on a 14-gallon tank of gas. You run out of gas on a deserted highway, but have 1/2 gallon of lawn mower gas with you. At the current rate of travel, how far will you be able to go and will you reach the nearest gas station 15 miles away?

OpenStudy (anonymous):

is i/2 gallon @Vianne

OpenStudy (anonymous):

1/2

OpenStudy (vianne):

Ok.. So, do you know how many miles per a gallon first?

OpenStudy (anonymous):

no

OpenStudy (vianne):

Ok, so we know that.. 380 = miles. 14 = gallon tank of gas. So if we do \[380 \div 14 = ?\] What is the answer?

OpenStudy (anonymous):

so if a divied 380/ 14 is 27.14

OpenStudy (vianne):

So we now have 27.14. But! You said we have .5 gallon tank of gas left.. But "27.14" is 1 gallon. We need to solve for .5 gallons. So.. We do \[27.14 \div .5 =\] What's the answer?

OpenStudy (anonymous):

5.428

OpenStudy (vianne):

No, not "5" but .5 or 0.5. Decimal not whole number :)

OpenStudy (anonymous):

ok

OpenStudy (vianne):

Found the answer yet? :D

OpenStudy (anonymous):

so is not C

OpenStudy (vianne):

Nope. That's for 1 gallon. You said you had 0.5 gallons left. We need to find how many miles are in 0.5 gallons.

OpenStudy (anonymous):

im soo confuse

OpenStudy (vianne):

380 / 14= 27.14(too many digits, so we put as 27.14) for now. 27.14 = miles per gallon 27.14/.5= ?=miles per half a gallon

OpenStudy (vianne):

If you can solve 27.14/.5 then you'll get your answer.

OpenStudy (vianne):

Still having trouble?

OpenStudy (vianne):

Well, I don't know what answer you picked. XD So I'll guess you picked D which is the correct answer. 380 / 14= 27.14(too many digits, so we put as 27.14) for now. 27.14 = miles per gallon 27.14/.5=13.57 13.57=miles per half a gallon

OpenStudy (vianne):

@rocky1718 :)